find the inverse of this matrix by using Gausian method


A= 4X + 5Y = 6


X + 5Y = 3


1
Expert's answer
2022-01-18T06:31:03-0500
"4X + 5Y = 6"

"X + 5Y = 3"

"A=\begin{bmatrix}
 4 & 5 \\
 1 & 5
\end{bmatrix}"

"\begin{bmatrix}
 4 & 5 & & 1 & 0 \\
 1 & 5 & & 0 & 1 
\end{bmatrix}"

"R_1=R_1\/4"


"\begin{bmatrix}
 1 & 5/4 & & 1/4 & 0 \\
 1 & 5 & & 0 & 1 
\end{bmatrix}"

"R_2=R_2-R_1"


"\begin{bmatrix}
 1 & 5/4 & & 1/4 & 0 \\
 0 & 15/4 & & -1/4 & 1 
\end{bmatrix}"

"R_2=4R_2\/15"


"\begin{bmatrix}
 1 & 5/4 & & 1/4 & 0 \\
 0 & 1 & & -1/15 & 4/15 
\end{bmatrix}"

"R_1=R_1-5R_2\/4"


"\begin{bmatrix}
 1 & 0 & & 1/3 & -1/3 \\
 0 & 1 & & -1/15 & 4/15 
\end{bmatrix}"

"A^{-1}=\begin{bmatrix}
 1/3 & -1/3 \\
 -1/15 & 4/15
\end{bmatrix}"